Object A and Object B are charged with +37.2 u coulombs and -68.1 μ coulombs, respectively Given Chat the wo objects are separated by 0.37 m. Which of the following statements is/are valid?




a. Object A attracts object B with a force of 20.5 N



b. Object 8 repels object A with a force of 166.4 N



c. Object A repels object B with a force of 166.4 N d. Object B attracts object A with a force of 166.4 N



e. Object B attracts object A with a force of 2015 N



f. Object A repels object 8 with a force of 20.5 N



g. Object B repels object A with a force of 20.5 N



HObject A attracts object 8 with a force of 166.4 N

1
Expert's answer
2022-03-13T18:52:09-0400

Force

"F=\frac{kq_1q_2}{r^2}"

Put value


"F=-\frac{9\times10^9\times37.2\times10^{-6}\times68.1\times10^{-6}}{0.37^2}=-166.4N"

Attractive force

Option (h) is correct option